# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v3

@@ II.3 Dynamical regimes
A Bautin bifurcation occurs when an unstable and a stable limit cycle collide and annihilate each other, a complexity also found in the Hodgkin-Huxley model Lee et al. (1998). The existence of a saddle-node of cycles suggests an unstable cycle nestled within a stable limit cycle. Therefore, in the tri-valued domain, initializing the system away from the fixed points is crucial to converge towards the limit cycle. Failure to do so results in convergence to the stable fixed points, leading to bistable behavior.
The provided paragraphs offer an overview of the key dynamical regimes and behaviors that one could expect in the local FHN model. Given the complex nature of these dynamics, a thorough exploration extends beyond the scope of this summary. For readers seeking an in-depth understanding, we suggest consulting Rocsoreanu’s work Rocsoreanu et al. (2012), specifically chapter 5, where these regimes are explored in greater detail.
@@ II.4 Time scale separation
Finally, we address the impact of the parameter $\varepsilon$ on the FHN model. Unlike $a$ and $b$, which influence the nullclines directly, $\varepsilon$ modifies the time scale separation, thereby adjusting the speed of dynamics in the variable $v$. This effect is evident in the vector field displayed in the first panel of Fig. 6A. It is important to note that in systems with pronounced time scale separation, the oscillation period is predominantly determined by the slower segments. As the dynamics in these segments accelerate, the oscillation period diminishes (Fig. 6A, right panel), leading to the disappearance of relaxation oscillations when the separation is minimal. Therefore, time scale separation, alongside multistability, plays a crucial role in enabling excitability and relaxation oscillations, which are central to the FHN model’s utility. Additionally, the transition to relaxation-like oscillations is significantly influenced by the degree of time scale separation. As depicted in Fig. 6B, increasing $\varepsilon$—which reduces time scale separation—results in oscillation amplitudes that vary with $a$, moving away from the constant amplitude characteristic of relaxation oscillations.
@@ III The diffusively coupled FitzHugh-Nagumo system
When developing the FHN model Nagumo et al. (1962), Nagumo et al. extended the original equation by incorporating diffusive coupling terms, facilitating the transmission of excitation pulses along the axon. This advancement was practically demonstrated through the construction of interconnected FHN circuits, which effectively mimicked pulse propagation, as depicted in Fig. 2 C. They also explored the thresholds at which these pulses become excited.
@@
Originally conceived for neuroscience applications, the FHN model FitzHugh (1961) and its spatial extension Nagumo et al. (1962) have been used beyond their initial scope, finding relevance across various disciplines. In cardiology, the model is celebrated for its ability to simulate spiral wave dynamics Franzone et al. (2006); Davidenko et al. (1995); Courtemanche et al. (1990); Keener (1988); Keener and Sneyd (2009); Aliev and Panfilov (1996); Biktashev (2002); Pertsov et al. (1993); Nash and Panfilov (2004); Watanabe et al. (2004); Sermesant et al. (2006); Cherubini et al. (2008); Glass and Josephson (1995); Le and Sotiropoulos (2013); Panfilov and Keener (1995); Nagai et al. (2000), while more generally in biology, it is recognized for generating traveling waves that effectively convey information Gelens et al. (2014); Noguchi (2023); Nolet et al. (2020a); Gambino et al. (2022, 2023); Cartwright et al. (1999); Duffy et al. (1998); Sneyd et al. (2017b); Lee (1997); Gelens et al. (2015). The model’s utility extends to chemistry and computational sciences Moayeri et al. (2020); Dehghan and Narimani (2018); Fuselier and Wright (2013); Rogers and McCulloch (1994); Kuramoto et al. (1998); Lin et al. (2004), where it aids in understanding complex systems. Beyond its interdisciplinary applications, the FHN model has been instrumental in investigating mathematical properties Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Keener (1986a); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Ohta et al. (1989); Iqbal et al. (2017); Elphick et al. (1990); Meron (1992); Petrich and Goldstein (1994); Hagberg and Meron (1994a); Goldstein et al. (1996); Vasiev (2004); Nekhamkina and Sheintuch (2006); Panfilov and Hogeweg (1993); Winfree (1991a); Zheng and Shen (2015); Tsyganov et al. (2014); Rombouts and Gelens (2020); Winfree (1991b); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Zaritski et al. (2004); Bär and Eiswirth (1993a); Kuznetsov et al. (2017); Shepelev et al. (2017); Browne et al. (2008); Guckenheimer and Kuehn (2010, 2009); Perc (2005b); Shima and Kuramoto (2004); Hakim and Karma (1999); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Métens et al. (1997); Zemskov and Epstein (2010); Hastings (1982); Lin et al. (2004); Bode (1997); Tobias and Knobloch (1998); Marée and Panfilov (1997); Shoji et al. (2007); Yang et al. (2003), establishing it as a fundamental model in theoretical studies. This section aims to provide a theoretical exploration into the existence and linear stability of stationary homogeneous solutions and the emergence of spatially structured dynamics such as traveling waves and extended patterns. For in-depth mathematical treatments, readers are directed to Cross and Greenside (2009).
@@ III.1 Linear stability analysis of stationary homogeneous solutions
Introducing spatial coupling into the analysis does not change the stationary spatially homogeneous solutions or their bifurcation diagrams. However, it necessitates the consideration of spatially extended perturbations in the stability analysis due to the potential influence of spatial feedback mechanisms. Such feedback, mediated by diffusion and nonlinearity, often becomes significant near bifurcation points where various effects balance out. Consequently, destabilization of the homogeneous solution may give rise to complex spatial patterns. To elucidate the nature of these instabilities, we examine small temporal and spatial perturbations around the equilibrium states, expressed as $(u,v)=(u_{h},v_{h})+\epsilon(\xi_{u},\xi_{v})e^{i{\bf k}\cdot{\bf x}+\sigma t}+c.c.$, where $\epsilon\ll 1$. This leads to an eigenvalue problem:
@@
$J\equiv\begin{pmatrix}1-D_{u}k^{2}-3u_{b}^{2}&-1\\ \varepsilon&-D_{v}k^{2}-\varepsilon b\end{pmatrix}.$ (21)
@@
${\rm Det}[J](k)=D_{u}D_{v}k^{4}+(3D_{v}u_{b}^{2}-D_{v}+D_{u}\varepsilon b)k^{2}\\ +\varepsilon(3bu_{b}^{2}-b+1),$ (22)
${\rm Tr}[J](k)=-(D_{u}+D_{v})k^{2}+1-3u_{b}^{2}-\varepsilon b.$ (23)
Notably, setting $k=0$ recovers the instabilities found in the non-spatial system, such as Hopf and saddle-node bifurcations. The dispersion relation $\sigma(k)$ reveals that a variety of coherent structures, including traveling waves and stationary spatial patterns, emerge from distinct instabilities. Linear stability analysis in the presence of spatial coupling for the FHN model unveils a complex array of dynamical regimes (for an in-depth discussion, see references Ei et al. (2008); Iqbal et al. (2017); Zheng and Shen (2015); Hagberg et al. (2006); Hagberg and Meron (1994b); Shepelev et al. (2017); Guckenheimer and Kuehn (2010, 2009); Gambino et al. (2023); Berezovskaya et al. (2008); Sneyd et al. (2017b); Tobias and Knobloch (1998)). The role of nonlinearity is paramount in determining the existence and stability of these predicted structures. Our focus here is on the most prominent spatially coherent structures supported by the FHN model.
@@ III.2 Turing patterns in one spatial dimension
Spatial patterns, resulting from self-organizing processes, are widely observed in nature and have been a subject of interest across various scientific fields. In developmental biology, for example, such patterns provide a mechanism for breaking the symmetry in the initially homogeneous tissue of an organism, guiding the development of complex structures Kondo and Miura (2010). Similarly, in ecology, the self-organization of vegetation patterns plays a crucial role in ecosystem resilience to environmental fluctuations Rietkerk et al. (2021). Despite the diverse manifestations of self-organization, the underlying mathematical frameworks share a foundational theory, first introduced by Alan Turing in 1952 Turing (1990). Turing’s theory of morphogenesis highlighted the essential elements for spatial self-organization within reaction-diffusion systems like the FHN model. In these systems, an activator substance ($u$) promotes the production of its inhibitor ($v$), which diffuses more rapidly and suppresses the activator in adjacent areas, establishing a feedback mechanism that generates regular spatial patterns, as depicted in Fig. 7. The FHN model, while primarily explored from a theoretical perspective, serves as a valuable tool for studying pattern formation, particularly given its oscillatory dynamics that contribute to the complexity of the observed patterns (see section III.6).
@@
Figure 8: Mapping of the Turing bifurcation in $(a,b)$ space. A. The phase diagram, highlighting Turing bifurcations T± (red), Hopf bifurcations H± (blue), and saddle-node bifurcation (black), with a zoomed-in view of the cusp (C) within the black rectangle. B. Bifurcation diagrams for selected $b$ values of $-0.5,0.5$, and $2$, as marked by vertical dashed lines in A, with parameters set to $D_{u}=1$, $D_{v}=5$, and $\varepsilon=0.01$.
@@
Pattern formation often emerges at the juncture of competing effects, typically near other bifurcations of a system’s homogeneous state where specific eigenvalues converge towards zero. In the FHN model, Turing instabilities Turing (1990), which lead to pattern formation, appear close to the saddle-node bifurcations, illustrated in Figs. 8 and 9. From a detailed perspective, a Turing or modulation instability marks the transition where a uniformly stable state becomes unstable to perturbations of a specific wavelength $k_{c}$. This critical point is identifiable in the dispersion relation when the growth rate of mode $k_{c}$ turns positive (Fig. 9B), signaling the temporal increase of this mode’s amplitude and the onset of spatially periodic Turing patterns (Fig. 7). The bifurcation conditions, ${\rm Det}[J](k)=0$ and $\partial_{k}{\rm Det}[J](k)=0$, must be simultaneously met at a non-zero wavenumber $k=k_{c}$, leading to the critical wavelength and defining the instability’s location:
$k_{c}=\pm\sqrt{\dfrac{-3D_{v}{u_{b}}^{2}+D_{v}-D_{u}\varepsilon b}{2D_{u}D_{v}}},$ (24)
$u_{T}=\pm\sqrt{\dfrac{1}{3D_{v}}\left(D_{v}+D_{u}\varepsilon b\pm 2\sqrt{D_{u}D_{v}\varepsilon}\right)},$ (25)
where the criterion $D_{v}>3D_{v}{u_{b}}^{2}+D_{u}\varepsilon b$ must be satisfied.
@@
Figure 9: Exploring pattern formation with reduced time-scale separation. A. The $(a,b)$-phase diagram’s transformation is illustrated, highlighting the shifts in Hopf (blue) and Turing (red) bifurcations with varying $\varepsilon$, given $D_{u}=1$ and $D_{v}=5$. B. Turing pattern emergence at $a=0.025$, $b=1.26$, and $\varepsilon=0.5$, as indicated by the vertical dashed line in A. The growth rate as a function of wavenumber $k$ is shown for $a=a_{T}$ (left), followed by the resulting Turing pattern at the most unstable wavelength (center), and a bifurcation diagram tracing the pattern’s maxima and minima across varying $a$ values (right), utilizing the same parameters as in Fig. 8.
@@
The criterion for instability requires that the inhibitor’s diffusion rate surpasses that of the activator ($D_{v}>D_{u}$), aligning with Murray’s findings for a general system Murray (2001). The Turing bifurcation’s dependence on parameters $a$ and $b$ mirrors the Hopf bifurcation’s parameter dependence, as depicted in Fig. 8A with red and blue lines, respectively. These Hopf and saddle-node bifurcations correspond to the spatially uniform state instabilities shown in Fig. 5B. As $b$ increases significantly, the instabilities converge towards the saddle-node (SN) bifurcations (black lines) of the uniform state, indicating the homogeneous stationary state’s stabilization (or destabilization) with increasing (or decreasing) $b$.
@@
Particularly intriguing dynamics emerge near the cusp bifurcation $C$, where a region exhibiting coexisting bifurcations transitions from bistable to monostable regimes, highlighted in the inset of Fig. 8A. Different $b$ values shift the system from monostable to bistable regimes, influencing the bifurcations’ positions. The vertical dashed lines in Fig. 8A mark the bifurcation diagrams in Fig. 8B for $b=-0.5$, $b=0.5$, and $b=2$, showing that the Turing instability, marked with a red dot, occurs after the uniform states become Hopf unstable, denoted by a blue dot. At this $\varepsilon$ value, static Turing patterns are absent due to the faster growth of the homogeneous mode compared to the nonzero wave number mode.
@@
To achieve robust pattern formation, increasing $\varepsilon$ is necessary. Fig. 9A illustrates how the $(a,b)$-phase diagram transforms with varying $\varepsilon$, where the oscillatory region between the Hopf lines H narrows, and at $\varepsilon=1$, it confines to a lobe touching $C$ at $(a,b)=(0,1)$. It is noteworthy from Eq. (24) that similar effects occur when increasing $D_{u}/D_{v}$ instead of $\varepsilon$, since in both the dependence goes as ±$\propto\pm\sqrt{x\pm\sqrt{x}}$. Nonetheless, it is crucial to remember that pattern formation is contingent on $D_{v}>D_{u}$.
Figure 10: Dynamics of localized states through the lens of various front behaviors. A. Two contrasting front types - a ”tame” front (left) and a ”wild” front (right), with the tame front parameters set at $(D_{u},D_{v},a,b,\varepsilon)=(1,1,0,2,0.01)$ and the wild front at $(0.49,1,0.01,1.1,1)$. B. Sketch of how a traveling front can be created. C.The commonly used piecewise linear approximation leverages the separation of time scales to analytically predict front properties. D. Kymographs and their corresponding profiles for traveling tame fronts, employing the same parameters as the tame front with added asymmetry by adjusting $a$ to 0.1 (left) and -0.1 (right). E. Formation of localized states via front interactions, with $a$ values set at $10^{-5}$ (left) and $10^{-3}$ (right) for the parameters $(D_{u},D_{v},b,\varepsilon)=(0.49,1,1.1,1)$.
@@
With an increased value of $\varepsilon$, the difference in time scales between the variables diminishes, enhancing scale-dependent feedback, resulting in the Turing instability manifesting prior to the Hopf bifurcation. By fixing $b$ at 1.26 and setting $\varepsilon$ to 0.5, as depicted by the vertical dashed line, the bifurcation diagram reveals the emergence of spatially periodic patterns due to a Turing instability, illustrated in Fig. 9B. Near the Turing bifurcation point, the growth rate for modulated perturbations on the uniform state turns positive, signaling the start of instability, as indicated in Fig. 9B (left). Beyond this threshold, the dominant wavenumber $k_{u}$ initiates the formation of regular patterns, shown in Fig. 9B (center). The evolution of pattern maxima and minima with $a$, obtained from direct numerical simulations, is plotted with a red continuous line in Fig. 9B (right). Given these parameters, the pattern coexists with homogeneous solutions across a broad $a$ range, suggesting a potential subcritical emergence from the Turing instability, warranting further examination for confirmation. Unlike other pattern-forming systems, the symmetry in the FHN model leads to identical stationary patterns beyond T±. This symmetry implies that mode $k_{c}$ may bifurcate subcritically from T+, stabilize upon folding at lower $a$ values, and persist until nearing a second fold at a symmetrical $a$ value, then transitioning to the second Turing T-, akin to other models Thiele et al. (2019); Parra-Rivas et al. (2023b). To our knowledge, this bifurcation analysis of 1D patterns in the FHN model is new and unreported in existing literature, highlighting the need for more in-depth exploration of spatial structure bifurcations, particularly the stability of modes bifurcating after the Turing bifurcation far from the onset or the associated localized states.

### v4

@@ II.3 Dynamical regimes
A saddle-node bifurcation of limit cycles occurs when an unstable and a stable limit cycle collide and annihilate each other, a complexity also found in the Hodgkin-Huxley model Lee et al. (1998). The existence of a SNLC suggests an unstable cycle nestled within a stable limit cycle. Therefore, in the tri-valued domain, initializing the system away from the fixed points is crucial to converge towards the limit cycle. Failure to do so results in convergence to the stable fixed points, leading to bistable behavior.
This analysis offers an overview of the key dynamical regimes and behaviors that one could expect in the local FHN model. Given the complex nature of these dynamics, a thorough exploration extends beyond the scope of this summary. For readers seeking an in-depth understanding, we suggest consulting Rocsoreanu’s work Rocsoreanu et al. (2012), specifically chapter 5, where these regimes are explored in greater detail.
@@ II.4 Time scale separation
Finally, we address the impact of the parameter $\varepsilon$ on the FHN model. Unlike $a$ and $b$, which influence the nullclines directly, $\varepsilon$ modifies the time scale separation, thereby adjusting the speed of dynamics in the variable $v$. This effect is evident in the vector field displayed in the first panel of Fig. 6A. It is important to note that in systems with pronounced time scale separation, the oscillation period is predominantly determined by the slower segments. As the dynamics in these segments accelerate, the oscillation period diminishes (Fig. 6A, right panel), leading to the disappearance of relaxation oscillations when the separation is minimal. Therefore, time scale separation, alongside multistability, plays a crucial role in enabling excitability and relaxation oscillations, which are central to the FHN model’s utility. Additionally, the transition to relaxation-like oscillations is significantly influenced by the degree of time scale separation. As depicted in Fig. 6B, increasing $\varepsilon$ – which reduces time scale separation – results in oscillation amplitudes that vary with $a$, moving away from the constant amplitude characteristic of relaxation oscillations.
@@ III The diffusively coupled FitzHugh-Nagumo system
When developing the FHN model Nagumo et al. (1962), Nagumo et al. extended the original equation by incorporating diffusive coupling terms, facilitating the transmission of excitation pulses along the axon. This was practically demonstrated through the construction of interconnected FHN electronic circuits, which effectively mimicked pulse propagation, as depicted in Fig. 2C. They also explored the thresholds at which these pulses get excited.
@@
Originally conceived for neuroscience applications, the FHN model FitzHugh (1961) and its spatial extension Nagumo et al. (1962) have been used beyond their initial scope, finding relevance across various disciplines. In cardiology, the model is celebrated for its ability to simulate spiral wave dynamics Franzone et al. (2006); Davidenko et al. (1995); Courtemanche et al. (1990); Keener (1988); Keener and Sneyd (2009); Aliev and Panfilov (1996); Biktashev (2002); Pertsov et al. (1993); Nash and Panfilov (2004); Watanabe et al. (2004); Sermesant et al. (2006); Cherubini et al. (2008); Glass and Josephson (1995); Le and Sotiropoulos (2013); Panfilov and Keener (1995); Nagai et al. (2000), while more generally in biology, it is recognized for generating traveling waves that effectively convey information Gelens et al. (2014); Noguchi (2023); Nolet et al. (2020a); Gambino et al. (2022, 2023); Cartwright et al. (1999); Duffy et al. (1998); Sneyd et al. (2017a); Lee (1997); Gelens et al. (2015). The model’s utility extends to chemistry and computational sciences Moayeri et al. (2020); Dehghan and Narimani (2018); Fuselier and Wright (2013); Rogers and McCulloch (1994); Kuramoto et al. (1998); Lin et al. (2004), where it aids in understanding complex systems. Beyond its interdisciplinary applications, the FHN model has been instrumental in investigating mathematical properties Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Keener (1986b); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Ohta et al. (1989); Iqbal et al. (2017); Elphick et al. (1990); Meron (1992); Petrich and Goldstein (1994); Hagberg and Meron (1994a); Goldstein et al. (1996); Vasiev (2004); Nekhamkina and Sheintuch (2006); Panfilov and Hogeweg (1993); Winfree (1991a); Zheng and Shen (2015); Tsyganov et al. (2014); Rombouts and Gelens (2020); Winfree (1991b); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Zaritski et al. (2004); Bär and Eiswirth (1993b); Kuznetsov et al. (2017); Shepelev et al. (2017); Browne et al. (2008); Guckenheimer and Kuehn (2010, 2009); Perc (2005a); Shima and Kuramoto (2004); Hakim and Karma (1999); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Métens et al. (1997); Zemskov and Epstein (2010); Hastings (1982); Lin et al. (2004); Bode (1997); Tobias and Knobloch (1998); Marée and Panfilov (1997); Shoji et al. (2007); Yang et al. (2003), establishing it as a fundamental model in theoretical studies. This section aims to provide a theoretical exploration into the existence and linear stability of stationary homogeneous solutions and the emergence of spatially structured dynamics such as traveling waves and extended patterns. For in-depth mathematical treatments, readers are directed to Cross and Greenside (2009).
@@ III.1 Linear stability analysis of stationary homogeneous solutions
Introducing spatial coupling into the analysis does not change the stationary spatially homogeneous solutions or their bifurcation diagrams. However, it requires the consideration of spatially extended perturbations in the stability analysis. Spatial feedback, mediated by diffusion and nonlinearity, often becomes significant near bifurcation points where various effects balance out. Consequently, destabilization of the homogeneous solution may give rise to complex spatial patterns. To elucidate the nature of these instabilities, we examine small temporal and spatial perturbations around the homogeneous states, expressed as $(u,v)=(u_{\text{HS}},v_{\text{HS}})+\epsilon(\xi_{u},\xi_{v})e^{i{\mathbf{k}}\cdot{\mathbf{x}}+\sigma t}+c.c$., where $\epsilon\ll 1$. This leads to an eigenvalue problem:
@@
$J\equiv\begin{pmatrix}1-D_{u}k^{2}-3u_{\text{HS}}^{2}&-1\\ \varepsilon&-D_{v}k^{2}-\varepsilon b\end{pmatrix}.$ (21)
@@
$\begin{array}[]{ll}{\rm Det}[J](k)=D_{u}D_{v}k^{4}+(3D_{v}u_{\text{HS}}^{2}-D_{v}+D_{u}\varepsilon b)k^{2}+\varepsilon(3bu_{\text{HS}}^{2}-b+1),\vspace{3pt}\\ {\rm Tr}[J](k)=-(D_{u}+D_{v})k^{2}+1-3u_{\text{HS}}^{2}-\varepsilon b.\end{array}$ (22)
Setting $k=0$ recovers the instabilities found in the non-spatial system, such as Hopf and saddle-node bifurcations. The dispersion relation $\sigma(k)$ reveals that a variety of coherent structures, including traveling waves and stationary spatial patterns, emerge from distinct instabilities (for an in-depth discussion, see Refs. Ei et al. (2008); Iqbal et al. (2017); Zheng and Shen (2015); Hagberg et al. (2006); Hagberg and Meron (1994b); Shepelev et al. (2017); Guckenheimer and Kuehn (2010, 2009); Gambino et al. (2023); Berezovskaya et al. (2008); Sneyd et al. (2017a); Tobias and Knobloch (1998)). The role of nonlinearity is critical in determining the existence and stability of these predicted structures. Our focus here is on the most prominent spatially coherent structures supported by the FHN model.
@@ III.2 Turing patterns in one spatial dimension
Spatial patterns, resulting from self-organizing processes, are widely observed in nature and have been a subject of interest across various scientific fields. In developmental biology, for example, such patterns provide a mechanism for breaking the symmetry in the initially homogeneous tissue of an organism, guiding the development of complex structures Kondo and Miura (2010). Similarly, in ecology, the self-organization of vegetation patterns plays a crucial role in ecosystem resilience to environmental fluctuations Rietkerk et al. (2021). Despite the diverse manifestations of self-organization, the underlying mathematical frameworks share a foundational theory, first introduced by Alan Turing in 1952 Turing (1990). Turing’s theory of morphogenesis highlighted the essential elements for spatial self-organization within reaction-diffusion systems like the FHN model. In these systems, an activator substance ($u$) promotes the production of its inhibitor ($v$), which diffuses more rapidly and suppresses the activator in adjacent areas, establishing a feedback mechanism that generates regular spatial patterns, as depicted in Fig. 7. The FHN model, while primarily explored from a theoretical perspective, serves as a valuable tool for studying pattern formation, particularly given its oscillatory dynamics that contribute to the complexity of the observed patterns (see III.6).
@@
Figure 8: Mapping of the Turing bifurcation in $(a,b)$ space. A. The phase diagram, highlighting Turing bifurcations T±, Hopf bifurcations H±, and saddle-node bifurcation, with a zoomed-in view of the cusp (C) within the black rectangle. B. Bifurcation diagrams for selected $b$ values of $-0.5,0.5$, and $2$, as marked by vertical dashed lines in A, with parameters set to $D_{u}=1$, $D_{v}=1$, and $\varepsilon=0.01$.
@@
Pattern formation often emerges at the juncture of competing effects, typically near other bifurcations of a system’s homogeneous state where specific eigenvalues converge towards zero. In the FHN model, Turing instabilities Turing (1990), which lead to pattern formation, appear close to the saddle-node bifurcations, illustrated in Figs. 8 and 9. A Turing instability, sometimes also called a modulation instability, marks the transition where a uniformly stable state becomes unstable to perturbations of a specific wavelength $k_{c}$. This critical point is identifiable in the dispersion relation when the growth rate of mode $k_{c}$ turns positive (Fig. 9B), signaling the temporal increase of this mode’s amplitude and the onset of spatially periodic Turing patterns (Fig. 7). The bifurcation conditions, ${\rm Det}[J](k)=0$ and $\partial_{k}{\rm Det}[J](k)=0$, must be simultaneously met at a non-zero wavenumber $k=k_{c}$, leading to the critical wavelength and defining the instability’s location:
$k_{c}=\pm\sqrt{\dfrac{-3D_{v}{u_{\text{HS}}}^{2}+D_{v}-D_{u}\varepsilon b}{2D_{u}D_{v}}},$ (23)
$u_{\text{T}}=\pm\sqrt{\dfrac{1}{3D_{v}}\left(D_{v}+D_{u}\varepsilon b\pm 2\sqrt{D_{u}D_{v}\varepsilon}\right)},$ (24)
where the criterion $D_{v}>3D_{v}u_{\text{HS}}^{2}+D_{u}\varepsilon b$ must be satisfied.
@@
Figure 9: Exploring pattern formation with reduced time-scale separation. A. The $(a,b)$-phase diagram’s transformation is illustrated, highlighting the shifts in Hopf and Turing bifurcations with varying $\varepsilon$, given $D_{u}=1$ and $D_{v}=5$. B. Turing pattern emergence at $a=0.025$, $b=1.26$, and $\varepsilon=0.5$, as indicated by the vertical dashed line in A. The growth rate as a function of wavenumber $k$ is shown for $a=a_{T}$ (left), followed by the resulting Turing pattern at the most unstable wavelength (center), and a bifurcation diagram tracing the pattern’s maxima and minima across varying $a$ values (right), utilizing the same parameters as in Fig. 8.
@@
The criterion for instability requires that the inhibitor’s diffusion rate surpasses that of the activator ($D_{v}>D_{u}$), aligning with Murray’s findings for a general system Murray (2001). The Turing bifurcation’s dependence on parameters $a$ and $b$ mirrors the Hopf bifurcation’s parameter dependence, as depicted in Fig. 8A. These Hopf and saddle-node bifurcations correspond to the spatially uniform state instabilities shown in Fig. 5B. As $b$ increases significantly, the Turing and Hopf instabilities converge towards the saddle-node (SN) bifurcations (black lines) of the uniform state, indicating the homogeneous stationary state’s stabilization (or destabilization) with increasing (or decreasing) $b$.
@@
Particularly intriguing dynamics emerge near the cusp bifurcation $C$, where a region exhibiting coexisting bifurcations transitions from bistable to monostable regimes, highlighted in the inset of Fig. 8A. Different $b$ values shift the system from monostable to bistable regimes, influencing the bifurcations’ positions. The vertical dashed lines in Fig. 8A mark the bifurcation diagrams in Fig. 8B for $b=-0.5$, $b=0.5$, and $b=2$, showing that the Turing instability, occurs after the uniform states become Hopf unstable. At this value of $\varepsilon$, static Turing patterns are absent due to the faster growth of the homogeneous mode compared to the nonzero wave number mode.
@@
To achieve robust pattern formation, increasing $\varepsilon$ is necessary. Fig. 9A illustrates how the $(a,b)$-phase diagram transforms with varying $\varepsilon$, where the oscillatory region between the Hopf lines H narrows, and at $\varepsilon=1$, it confines to a lobe touching $C$ at $(a,b)=(0,1)$. It is noteworthy from Eq. 23 that similar effects occur when increasing $D_{u}/D_{v}$ instead of $\varepsilon$, since in both the dependence goes as ±$\propto\pm\sqrt{x\pm\sqrt{x}}$. Nonetheless, it is crucial to remember that pattern formation is contingent on $D_{v}>D_{u}$.
@@
With an increased value of $\varepsilon$, the difference in time scales between the variables diminishes, enhancing scale-dependent feedback, resulting in the Turing instability manifesting prior to the Hopf bifurcation. By fixing $b$ at 1.26 and setting $\varepsilon$ to 0.5 (see the vertical dashed line in Fig. 9A), the bifurcation diagram reveals the emergence of spatially periodic patterns due to a Turing instability, illustrated in Fig. 9B. Near the Turing bifurcation point, the growth rate for modulated perturbations on the uniform state turns positive, signaling the start of instability, as indicated in Fig. 9B (left). Beyond this threshold, the dominant wavenumber $k_{u}$ initiates the formation of regular patterns, shown in Fig. 9B (center). The evolution of pattern maxima and minima with $a$, obtained from direct numerical simulations, is plotted with a continuous line in Fig. 9B (right). Given these parameters, the pattern coexists with homogeneous solutions across a broad range in the parameter $a$, suggesting a potential subcritical emergence from the Turing instability, warranting further examination for confirmation. Unlike other pattern-forming systems, the symmetry in the FHN model leads to identical stationary patterns beyond T±. This symmetry implies that mode $k_{c}$ may bifurcate subcritically from T+, stabilize upon folding at lower $a$ values, and persist until nearing a second fold at a symmetrical $a$ value, then transitioning to the second Turing T-, akin to other models Thiele et al. (2019); Parra-Rivas et al. (2023b). To our knowledge, this bifurcation analysis of 1D patterns in the FHN model is new and unreported in existing literature, highlighting the need for more in-depth exploration of spatial structure bifurcations, particularly the stability of modes bifurcating after the Turing bifurcation far from the onset or the associated localized states.